ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

19. History of Mathematics

1991 MS Classification: 01A99 (01A55, 01A60) Ion, Patrick D. F., Mathematical Reviews, Ann Arbor, MI 48104 USA The symbolism of mathematics: from pasigraphy in 1897 to MathML in 1998. In 1897 at the first International Congress of Mathematicians in Zürich, the logician E. Schröder delivered one of the few addresses: “Über Pasigraphie, ihren gegenwärtigen Stand und die pasigraphische Bewegung in Italien”. He disagreed with the distinguished chairman G. Peano, by thinking that Leibniz’s dream of a characteristica universalis had not yet been achieved, but he was sure that pasigraphy would be represented at all subsequent ICM. We still develop symbolism and notation with the intention of streamlining communication and facilitating calculation, as the new World Wide Web Consortium recommendation MathML (Mathematical Markup Language: http://www.w3.org/Math ) shows. The presentation sketches some of the evolution of notation during this period.